# Large Human Language Models: A Need and the Challenges

### v2

@@ 2.1 Motivation
Piantadosi et al. (1988) describe a fallacy in statistical models of the world pertaining to modeling individual observations that are part of a group, as if they are independent, a so-called ecological fallacy:
"Serious errors can result when an investigator makes the seemingly natural assumption that the inference from an ecological analysis must pertain either to individuals within the group or to individuals across groups."
Current LLMs exhibit a form of this ecological fallacy, whereby text sequences written by an author are treated as if independent and miss the opportunity to capture dependence Soni et al. (2022).
@@
This human language modeling problem generalizes the regular language modeling problem of predicting the next word conditioned on the previous words in a text sequence $X$ to also condition on a human context $H$. To train LLMs for this human language modeling problem, we need methods to both represent the human context and to include it in our training objective.
@@ 3.3 Challenges and Possible Solutions
C1: Modeling data and representational disparities. To capture the rich human context, we need access to datasets that provide relevant information covering users who are representative of the broad and diverse population (Piantadosi et al., 1988; Henrich et al., 2010; Johnson et al., 2022). Specifically, the challenges lie in obtaining datasets: (1) that provide access to user identifiers and historical language which allow us to differentiate the human source of the language, and associate explicit human attributes such as sociodemographic or personality attributes, (2) that do not amplify representational disparities (Shah et al., 2020) and span multiple domains such as healthcare (Bean et al., 2023), customer service (Adam et al., 2021), and education (Klein and Nabi, 2019).
@@ Acknowledgments
This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #585968, and a grant from the CDC/NIOSH (U01 OH012476). The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, any other government organization, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.

### v3

@@ 2.1 Motivation
Robinson (1950) describes a fallacy in statistical models of the world pertaining to modeling individual observations that are part of a group, as if they are independent, a so-called ecological fallacy. Current LLMs exhibit a form of this ecological fallacy, whereby text sequences written by an author are treated as if independent and miss the opportunity to capture dependence Soni et al. (2022). Conversely, in the absence of a notion of authors, the LLMs can be seen as modeling the documents from many different people as if they were generated by a single universal author.
@@
This human language modeling problem generalizes the regular language modeling problem of predicting the next word conditioned on the previous words in a text sequence $X$ to also condition on a human context $H$ [^1] This framing can also be formulated for non-autoregressive language modeling.. To train LLMs for this human language modeling problem, we need methods to both represent the human context and to include it in our training objective.
@@ 3.3 Challenges and Possible Solutions
C1: Modeling data and representational disparities. To capture the rich human context, we need access to datasets that provide relevant information covering users who are representative of the broad and diverse population (Henrich et al., 2010; Johnson et al., 2022). Specifically, the challenges lie in obtaining datasets: (1) that provide access to user identifiers and historical language which allow us to differentiate the human source of the language, and associate explicit human attributes such as sociodemographic or personality attributes, (2) that do not amplify representational disparities (Shah et al., 2020) and span multiple domains such as healthcare (Bean et al., 2023), customer service (Adam et al., 2021), and education (Klein and Nabi, 2019).
@@ Acknowledgments
This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200005, and a grant from the CDC/NIOSH (U01 OH012476). The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, any other government organization, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.